OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 4, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = (2 points)

OpenStudy (anonymous):

@alassen

OpenStudy (anonymous):

out of curiosity, is it supposed to be f(n) = f(1) + f(2) + ... + f(n-1)? If so then you would just sum all preceding terms like so: f(3) = f(1) + f(2) = 2 + 4 =6 f(4) = f(1) + f(2) + f(3) = 2 + 4 + 6 = 12 f(5) = f(1) + f(2) + f(3) + f(4) = 2+4+6+12 =24 that's just the way I interpreted it first. if it is how you wrote it, ( f(n) = f(1) + f(2) + f(n - 1)) then you would still have to work your way up but the numbers would be slightly different: f(3) = f(1) + f(2) + f(n-1) = 2 + 4 + f(3-1) =6 + 4 = 10 f(4) = f(1) + f(2) + f(4-1) = 2 + 4 + f(3) = 6+10 = 16 f(5) = f(1) + f(2) + f(5-1) = 2 + 4 + f(4) = 6+16 = 22